REGISTRATION FORM - ISUME2

REGISTRATION DETAILS: Name and Title: Institution: Address for correspondence:

Email Address: Telephone Number: DISCUSSION GROUP: Please indicate which Discussion Group you would like to join: CONFERENCE DINNER: The conference dinner, which will be in Ouzo’s of Baggot Street on Friday, 10 December, will cost 35 . I will be attending the conference dinner: YES NO REGISTRATION FEE: Registration fee: 25 For students (undergraduate or postgraduate) who wish to attend, there is limited funding available to pay the 25 registration fee. Are you eligible for this waiver? YES NO Payment by cheque or bankdraft - to be made payable to ISUME-UCD Please send registration form and cheque for registration and/or conference dinner to: Dr Maria Meehan, Department of Mathematics, Science Lecture Building, University College Dublin, Belfield, Dublin 4, Ireland In order to assist with the organisation of ISUME2, please register before Friday, 3 December, 2004.